# On energy and its positivity in spacetimes with an expanding flat de Sitter background

### v1

@@ 1. Introduction
Let $(\mathscr{M},\mathscr{g})$ be a four-dimensional spacetime, i.e., a time-oriented Lorentzian manifold. An initial data set is a quadruple $(M,g,k,\lambda)$, where $M$ is a spacelike hypersurface, $g$ is the metric induced on $M$, $k$ is the second fundamental form of $M$ in $\mathscr{M}$, and $\lambda$ is a constant. We note that the pair $(g,k)$ may depend on $\lambda$. In what follows we assume that $\lambda\geq 0.$ When $\lambda=0$, we adopt the classical notation $(M,g,k)$. In particular, if $u$ is the future-directed unit timelike vector field, then, in our convention,
@@
where $X,Y$ are tangent vector fields on $M$ and $\nabla$ is the spacetime Levi-Civita connection. With this convention, a positive mean curvature of the slice corresponds to expansion towards the future.
The initial data set must satisfy the Einstein constraint equations, which arise from the Gauss–Codazzi decomposition of the Einstein field equations. In the presence of a cosmological constant $\Lambda=3\lambda^{2}$, these constraints take the form:
$2\rho+2\Lambda$ $=R_{g}+({\rm tr}_{g}k)^{2}-|k|_{g}^{2},$ (2)
@@
where $R_{g}$ is the scalar curvature of the induced metric $g$, $\rho$ denotes the energy density, and $J$ is the momentum density of the matter fields.
An initial data set is said to satisfy the dominant energy condition with respect to $\Lambda$ if the inequality
@@
holds.
We remark that the dominant energy condition used here is derived from the spacetime dominant energy condition applied to the stress-energy tensor $\mathscr{T}=\mathcal{G}+\Lambda\mathscr{g},$ where $\mathcal{G}$ is the Einstein tensor associated with $\mathscr{g}$. Specifically, we require that for every future-directed timelike vector $\mathscr{v}$, the vector $\mathscr{T}(\mathscr{v},\cdot)^{\sharp}$ is also future-directed timelike. Consequently, this dominant energy condition depends explicitly on $\Lambda$. This condition ensures that energy density dominates momentum flux, reflecting the physical requirement that energy flows cannot exceed the speed of light.
@@
Noether’s theorem (see [Arn89]), loosely speaking, states that conserved quantities correspond to symmetries of the system. Therefore, to define conserved quantities such as energy in general relativity, we look for global symmetries of the spacetime. These symmetries are generated by Killing vector fields, which are vector fields that preserve the metric along their flow (i.e. they generate isometries of the spacetime). In particular, a timelike Killing vector field corresponds to a time-translation symmetry and, by Noether’s theorem, gives rise to a conserved energy. However, most spacetimes do not possess such global symmetries. This is why asymptotically Minkowskian spacetimes are of particular interest: in these spacetimes the geometry approaches Minkowski space at infinity, where a timelike Killing vector field exists. This asymptotic symmetry provides the structure needed to define a conserved energy at infinity.
The ADM energy [ADM61] $E_{\mathrm{ADM}}$ plays a central role in the analysis of asymptotically flat spacetimes. It is derived from the Hamiltonian formulation of general relativity under the assumption of a fixed Minkowski background. It provides a notion of total energy for an isolated gravitational system and is invariant under asymptotically flat coordinate transformations [Bar86, Chr88]. This invariance guarantees that the energy is well-defined independently of the observer’s coordinate choice at infinity.
The ADM energy is defined for initial data sets $(M,g,k,\lambda=0)$ that satisfy specific fall-off conditions at spatial infinity, namely:
@@
where $\tau>\frac{1}{2}$. These asymptotic conditions ensure that the metric $g$ approaches the flat Euclidean metric and that the extrinsic curvature $k$ decays appropriately. Assuming integrability of $\mu$ and $J$, the energy is then defined as
@@
where $d\mu_{\delta}$ denotes the area measure on the coordinate sphere $S_{r}=\{|\mathbf{x}|=r\}$ induced by the Euclidean metric $\delta$. We remark that in this setting also the linear momentum and ADM mass would be well defined.
A fundamental result in mathematical relativity is the Positive Energy Theorem. It was first proved for the time symmetric case $k\equiv 0$ independently by Schoen and Yau [SY79] and Witten [Wit81]. It was later extended for the case of a general $k$ with appropriate asymptotic decay by Schoen and Yau [SY81a, SY81b]. The theorem states that, under the assumption that the dominant energy condition holds, the ADM energy is non-negative:
@@
Equality occurs if and only if the initial data $(M,g,k)$ arises from a spacelike slice of flat Minkowski spacetime.
@@
When $\Lambda>0$, the situation becomes more delicate. A Positive Energy Theorem in this setting has been established by Borghini and Mazzieri [BM18, BM20] for compact static metrics. We also note the connection with rigidity and the Min–Oo conjecture (see [BMN11], where it was shown to be false). In contrast with [BM18, BM20], our focus is on umbilic initial data sets, which are not static.
The question we want to address is how to describe isolated gravitational systems within an expanding universe. Observational evidence of the accelerated expansion of the universe strongly supports a cosmological model with a positive cosmological constant $\Lambda>0$.
To study isolated systems in this framework, we adopt a background spacetime model that reflects this expansion. The flat expanding de Sitter model provides a natural choice, as it captures the large-scale behavior of a universe dominated by $\Lambda$-vacuum energy while retaining the mathematical simplicity of spatial flatness. Our goal is to find a definition of energy in this setting and establish a corresponding positivity result.
@@
Recall that the De Sitter spacetime is a maximally symmetric solution to Einstein’s equations with a positive cosmological constant $\Lambda>0$. It serves as a model for a universe undergoing accelerated expansion, which is supported by current observations. In comoving coordinates, the metric for the de Sitter spacetime takes the form
@@
where
@@
Therefore, this particular foliation of de Sitter spacetime yields an ideal background to model spatially flat systems which undergo cosmological expansion.
@@
In the quasi-local setting, one considers an initial data set $(\Omega,g,k,\lambda)$, where $\Omega$ is compact with some boundary component [^1] In general, $\Omega$ can have multiple boundary components. The perspective here is that $\Sigma$ is an outer boundary component that encompasses multiple inner boundary components that could represent black hole horizons. $\Sigma:=\partial\Omega$ of positive Gauss curvature. By the Weyl embedding theorem [Nir53, Pog52], there is an isometric embedding of $\Sigma$ within Euclidean space $\mathbb{R}^{3}$ which bounds a compact and convex subset $\Omega_{0}$. Let $H_{0}$ and $H$ denote the mean curvatures of $\Sigma$ within $\mathbb{R}^{3}$ and $\Omega$, respectively. (In our convention, the mean curvature is the trace of the second fundamental form with respect to the outward unit normal.) Then we define the energy $E_{\lambda}$ as a direct generalization of the Liu–Yau energy [LY03, LY06]:
@@
As for the Liu–Yau energy, $E_{\lambda}$ can be expressed as the difference between the norms of the codimension-two mean curvature vectors in the relevant spacetimes. Indeed it’s not hard to see that
@@
Here $\vec{H}$ is the mean curvature vector of $\Sigma$ (assumed to be spacelike) within the original spacetime $(\mathscr{M},\mathscr{g})$ containing $(\Omega,g,k,\lambda)$, while $\vec{H}_{0}$ is the mean curvature vector of $\Sigma$ within the de Sitter spacetime $(\mathscr{M}_{0},\mathscr{g}_{0})$ with $\Lambda=3\lambda^{2}$ containing the compact initial data set $(\Omega_{0},\delta,k_{0}=\lambda\delta,\lambda)$ within a constant $t$ slice. Both mean curvature vectors are spacelike in this setting. See figure 2 for a comparison between the Liu-Yau energy and $E_{\lambda}$.
@@
The structure of the paper is as follows. In Section 2 we state the main conjecture and our main results, Theorem 2.3 and Corollary 2.4. Section 3 reviews the relevant notions of quasi-local energy developed in a Minkowskian background, together with the associated positivity results. In Section 4 we provide the relevant lemmas and prove our main results. Section 5 explores examples of our results. Section 5.1 reviews the Schwarzschild–de Sitter spacetime in flat expanding coordinates and computes the quantities appearing in Theorem 2.3. Section 5.2 constructs many umbilic initial data sets $(\Omega,g,k=\lambda g,\lambda)$ satisfying the $\Lambda$-vacuum constraint equations (2) where our results apply. Finally, in Section 6 we summarize our findings and discuss future directions.
@@
Figure 2. Left: the Penrose diagram of Minkowski spacetime. To compute the Liu–Yau energy, the surface $\Sigma$ is isometrically embedded in a constant $t$-slice of Minkowski spacetime. Right: the Penrose diagram of de Sitter spacetime. The surface is also isometrically embedded in a constant $t$-slice of the flat de Sitter spacetime and lies within the cosmological horizon (shown in red).
@@ 2. Main conjecture and theorem
Consider an initial data set $(\Omega,g,k,\lambda)$ where $\Omega$ is compact with some boundary component $\Sigma=\partial\Omega$ and $\lambda=\sqrt{\Lambda/3}\geq 0$, where $\Lambda$ is the fixed cosmological constant of the background de Sitter space. Assume that $\Sigma$ is a closed, 2-dimensional surface of positive Gauss curvature. By the Weyl embedding theorem, which was solved independently by Nirenberg and Pogorelov [Nir53, Pog52], $\Sigma$ can be isometrically embedded into the Euclidean space $\mathbb{R}^{3}$ with the image of $\Sigma$ as the boundary of a compact and convex domain $\Omega_{0}\subset\mathbb{R}^{3}$; moreover, this embedding is unique up to Euclidean rigid motions. Let $H_{0}$ be the mean curvature of the embedding in $\mathbb{R}^{3}$ with respect to the outward normal. In the initial data $\Omega$, let $H$ be the mean curvature of $\Sigma$ with respect to outward normal.
Our first conjecture states that the energy $E_{\lambda}$ should be nonnegative whenever the initial data set satisfies the dominant energy condition with respect to $\Lambda$ and the Weyl embedding of $\Sigma$ in $\mathbb{R}^{3}$ lies within the cosmological horizon; as in Liu-Yau [LY03, LY06], we also require that the mean curvature vectors of $\Sigma$ are spacelike within $\Omega$ and the Weyl embedding in $\mathbb{R}^{3}$ so that $E_{\lambda}$ is well defined.
@@
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
@@
In the perfectly umbilic setting, $k=\lambda g$, the question becomes purely Riemannian. Indeed, from the first constraint equation, the dominant energy condition with $\Lambda=3\lambda^{2}$ is equivalent to requiring nonnegative of scalar curvature, $R_{g}\geq 0$. This leads to the purely Riemannian version of Conjecture 2.1.
@@
Let $(\Omega,g,k=\lambda g,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
@@
Addressing these conjectures involve two challenges. The first is geometric in nature: one would like to determine under which conditions the surface $\Sigma$ can be isometrically embedded inside the cosmological horizon. The second concerns the positivity of the energy $E_{\lambda}$ itself. We will show that both of these concerns can be resolved for some values of $\lambda$.
We start by introducing a constant $\alpha_{\lambda}=\alpha(g,k,\lambda)$ as:
$\alpha_{\lambda}=\begin{cases}\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}&\text{if }\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\\ \min\left\{\sqrt{\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right)},\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}\right\}&\text{if }\sqrt{K_{\min}}\geq\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\end{cases}.$ (19)
In this expression, $K_{\min}$ is the minimum of the Gauss curvature $K$ on $\Sigma$, $r_{\Sigma}$ is the area radius of $\Sigma$, and $E_{\rm LY}$ is the Liu-Yau energy of the initial data (see equation (24) in the next section). We note that the definition of $\alpha_{\lambda}$ does not involve $\lambda$ explicitly; however, an implicit dependence may occur if either the metric $g$ or the second fundamental form $k$ (and hence the Liu–Yau energy $E_{\rm LY}$) depends on $\lambda$.
@@
Let $(\Omega,g,k,\lambda)$ for $\lambda\geq 0$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}$ holds. If $\lambda\leq\alpha_{\lambda}$, where $\alpha_{\lambda}$ is given by (19), then the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and the energy (15) is nonnegative:
@@
Moreover, if $0<\lambda<\alpha_{\lambda}$, then $E_{\lambda}>0$; if $\lambda=\alpha_{\lambda}>0$ and $E_{\lambda}=0$, then $\Sigma$ is isometric to a round sphere.
Remark.
The condition $\lambda\leq\alpha_{\lambda}$ may be always satisfied in certain settings; an explicit example is provided by the Schwarzschild–de Sitter case discussed in Section 5.1.
@@
    Fixed geometric data. One may fix an initial data set $(\Omega,g,k)$ that is completely independent of $\lambda$. In this case, $\lambda$ acts as an external parameter, and $\alpha_{\lambda}$ depends only on the intrinsic geometry of $\Sigma$ and on the Liu–Yau energy $E_{\rm LY}$ computed from the fixed data and hence is independent of $\lambda$. Therefore the theorem provides an interval $0,\lambda_{\max}]$, where $\lambda_{\max}=\alpha$, of admissible values of $\lambda$ for which the data embeds within the cosmological horizon and the quasi-local energy $E_{\lambda}$ is nonnegative. In this case, the rigidity part of Theorem 2.3 can be understood as if $\lambda=\lambda_{\rm max}>0$ and $E_{\lambda}=0$, then $\Sigma$ is isometric to a round sphere.
@@
    $\lambda$-dependent data. Alternatively, one may consider a $\lambda$-dependent family of data $(\Omega,g(\lambda),k(\lambda))$, for instance determined by the constraint equations or by prescribing an umbilic second fundamental form. In this case, both $E_{\rm LY}$ and $\alpha_{\lambda}$ vary with $\lambda$, and the theorem can be applied pointwise by verifying the condition $\lambda\leq\alpha_{\lambda}$ for each value of $\lambda$.

### v2

@@ 1. Introduction
The question we want to address is how to describe energy in isolated gravitational systems within an expanding universe. In physics, isolated gravitational systems are meant to represent localized matter sources which generate a gravitational field decaying as one moves further away from these sources. Traditionally, such systems have been described in terms of spacetimes which, in some meaningful sense, asymptote to Minkowski space $\mathbb{R}^{3,1}$ as one moves to space-like infinity. This can be captured by considering solutions to the Einstein equations arising by the evolution of initial data which asymptote to the canonical initial data for $\mathbb{R}^{3,1}$, i.e., a totally geodesic spacelike hyperplane, as we move to space-like infinity. The physical intuition behind this choice arises by considering $\mathbb{R}^{3,1}$ as the vacuum solution to the Einstein equations that one would approach by moving continuously to exceedingly weaker gravitational fields created by a localized matter distribution. Nevertheless, observational evidence of the accelerated expansion of the universe strongly supports a cosmological model with a positive cosmological constant $\Lambda>0$, where the corresponding maximally symmetric cosmological $\Lambda$-vacuum solution to the Einstein equations is not given by $\mathbb{M}^{4}$, but by de Sitter space. Thus, as we move further away from a localized energy density, one should conjecture that the corresponding spacetime should approach a de Sitter background solution. In this paper, taking this point of view, we aim to present a first candidate for energy in this expanding setting and prove its positivity in some scenarios.
Let us highlight that, in the traditional viewpoint concerning asymptotically Mikowskian systems as models for an isolated system, asymptotic symmetries of $\mathbb{R}^{3,1}$ provide natural tools to propose a notion of energy attached to such a system. This is because, although generic spacetimes possess no symmetries, the timelike Killing fields in the asymptotic background of $\mathbb{R}^{3,1}$ can be used to produce such quantities through well-known methods [CHO84, CHR12, WZ00]. In the context of de Sitter space, there are no global timelike Killing fields, producing a challenge if one intendeds to follow a similar approach and identify a conserved quantity canonically related to a time-translational background symmetry. Moreover, it is not clear that within an expanding universe with fixed cosmological constant one should expect such a conservation law: On the one hand, matter should get increasingly diluted while on the other hand any contribution from $\Lambda$ remains constant as a density while space expands. These and other challenges motivate us to propose instead a quasi-local notion of energy, attaching such a notion to bounded subsets within the cosmological horizon of de Sitter space, instead of a global one, which may be typically guided by such conservation principles. We take this to be a promising first avenue in a program related to better defining isolated systems and their energy in expanding cosmological spacetimes.
Given the above comments, we point out that it might be an interesting related question to understand if a natural global notion of energy could be introduced by considering the background global conformal timelike Killing fields of de Sitter space. In particular it is of interest to investigate whether conformal Killing fields could introduce monotonicity laws addressing the subtleties outlined above. Pursuing this alternative is outside the scope of our current work, and inspiration could be drawn from [CHO84, ALM25]. If such a program could be successfully carried out, it would then be interesting to understand its relation to our quasi-local definition we put forth.
The structure of the paper is as follows. In Section 2 we define our quasi-local notion of energy with an expanding flat de Sitter background, see Definition 2.1. In Section 3 we state the main conjecture and our main results, Theorem 3.3 and Corollary 3.4. Section 4 reviews the relevant notions of quasi-local energy developed in a Minkowskian background, together with the associated positivity results. In Section 5 we provide the relevant lemmas and prove our main results. Section 6 explores examples of our results. Section 6.1 reviews the Schwarzschild–de Sitter space in flat expanding coordinates and computes the quantities appearing in Theorem 3.3. Section 6.2 uses a blow-up argument of Schoen to construct many umbilic initial data sets satisfying the $\Lambda$-vacuum constraint equations (2) where our results apply. Finally, in Section 7 we summarize our findings and discuss future directions.
2. Defining energy with an expanding flat de Sitter background
An initial data set is a triple $(M,g,k)$, where $(M,g)$ is a Riemannian 3-manifold and $k$ is a symmetric $(0,2)$ tensor on $M$. Two-sided spacelike hypersurfaces in a spacetime $(\mathscr{M},\mathscr{g})$ are examples of initial data sets; in this case $k$ is the second fundamental form of $M$ within $\mathscr{M}$. In particular, if $u$ is the future-directed unit timelike vector field, then, in our convention,
@@ 2. Defining energy with an expanding flat de Sitter background
where $X,Y$ are tangent vector fields on $M$ and $\nabla$ is the spacetime Levi-Civita connection. With this convention, positive mean curvature of the initial data set corresponds to positive expansion towards the future.
Given $\Lambda\in\mathbb{R}$ (the cosmological constant), we define the energy density $\rho$ and momentum density $J$ via,
$\rho+\Lambda$ $=\frac{1}{2}\bigg(R_{g}+({\rm tr}_{g}k)^{2}-|k|_{g}^{2}\bigg),$ (2)
@@
where $R_{g}$ is the scalar curvature of the induced metric $g$. We refer to these equations as the constraint equations; they arise from the Gauss–Codazzi decomposition of the Einstein field equations.
An initial data set satisfies the dominant energy condition with respect to $\Lambda$ if the inequality
@@
holds. We remark that the dominant energy condition used here is derived from the spacetime dominant energy condition applied to the stress-energy tensor $\mathscr{T}=\mathcal{G}+\Lambda\mathscr{g},$ where $\mathcal{G}$ is the Einstein tensor associated with the spacetime metric $\mathscr{g}$. Specifically, we require that for every future-directed timelike vector $\mathscr{v}$, the vector $\mathscr{T}(\mathscr{v},\cdot)^{\sharp}$ is also future-directed timelike. Consequently, this dominant energy condition depends explicitly on $\Lambda$. This condition ensures that energy density dominates momentum flux, reflecting the physical requirement that energy-momentum flows do not exceed the speed of light.
@@
Noether’s theorem (see [ARN89]), loosely speaking, states that conserved quantities correspond to symmetries of the system. Therefore, to define conserved quantities such as energy in general relativity, we look for global symmetries of the spacetime. These symmetries are generated by Killing vector fields, which are vector fields that preserve the metric along their flow (i.e. they generate isometries of the spacetime). In particular, a timelike Killing vector field corresponds to a time-translation symmetry and, by Noether’s theorem, gives rise to a conserved energy. However, general spacetimes do not possess such global symmetries. This is why asymptotically Minkowskian spacetimes are of particular interest: in these spacetimes the geometry approaches Minkowski space at infinity, where a timelike Killing vector field exists. This asymptotic symmetry provides the structure needed to define a conserved energy at infinity.
The ADM energy [ADM61] $E_{\mathrm{ADM}}$ plays a central role in the analysis of asymptotically flat spacetimes. It is derived from the Hamiltonian formulation of general relativity under the assumption of a fixed Minkowski space background. It provides a notion of total energy for an isolated gravitational system and is invariant under asymptotically flat coordinate transformations [BAR86, CHR88]. This invariance guarantees that the energy is well-defined independently of the observer’s coordinate choice at infinity.
The ADM energy is defined for initial data sets $(M,g,k)$ that satisfy specific fall-off conditions at spatial infinity, namely:
@@
where $\tau>\frac{1}{2}$. These asymptotic conditions ensure that the metric $g$ approaches the flat Euclidean metric $\delta$ and that the extrinsic curvature $k$ decays appropriately. Assuming integrability of $\mu$ and $J$, the energy is defined as
@@
where $d\mu_{\delta}$ denotes the area measure on the coordinate sphere $S_{r}=\{|\mathbf{x}|=r\}$ induced by the Euclidean metric $\delta$. We remark that in this setting the linear momentum and ADM mass are also be well defined.
A fundamental result in mathematical relativity is the Positive Energy Theorem. It was first proved for the time symmetric case $k\equiv 0$ independently by Schoen and Yau [SY79] and Witten [WIT81]. It was later extended for the case of a general $k$ with appropriate asymptotic decay by Schoen and Yau [SY81b, SY81a]. The theorem states that, under the assumption that the dominant energy condition holds, the ADM energy is non-negative:
@@
Equality occurs if and only if the initial data set $(M,g,k)$ arises from a spacelike slice of flat Minkowski spacetime.
@@
When $\Lambda>0$, the situation becomes more delicate. A Positive Energy Theorem in this setting has been established by Borghini and Mazzieri [BM18, BM20] for compact static metrics. See also [ABM26] for recent developments on mass-type invariants in the presence of a positive cosmological constant. We also note the connection with rigidity and the Min–Oo conjecture (see [BMN11], where it was shown to be false). In contrast with [BM18, BM20, ABM26], our focus is on initial data sets which expand, and hence are not static.
To study isolated systems in this framework, we adopt a background spacetime model that reflects the observed expansion of the universe. The flat expanding de Sitter model provides a natural choice, as it captures the large-scale behavior of a universe dominated by $\Lambda$-vacuum energy while retaining the mathematical simplicity of spatial flatness. Our goal is to find a definition of energy in this setting and establish a corresponding positivity result.
@@
Recall that de Sitter space is the maximally symmetric solution to Einstein’s equations with a positive cosmological constant $\Lambda>0$. It serves as a model for a universe undergoing accelerated expansion, which is supported by current observations. In comoving coordinates, the metric for the de Sitter space takes the form
@@
where [^1] The relation $\Lambda=3\lambda^{2}$ will hold throughout the rest of the paper.
@@
Therefore, this particular foliation of de Sitter space yields an ideal background to model spatially flat systems which undergo cosmological expansion.
@@
In the quasi-local setting, one considers an initial data set $(\Omega,g,k)$, where $\Omega$ is compact with with boundary [^2] In general, $\Omega$ can have multiple boundary components. $\Sigma:=\partial\Omega$ of positive Gauss curvature. By the Weyl embedding theorem [NIR53, POG52], there is an isometric embedding of $\Sigma$ within Euclidean space $\mathbb{R}^{3}$ which bounds a compact and convex subset $\Omega_{0}$. Let $H_{0}$ and $H$ denote the mean curvatures of $\Sigma$ within $\mathbb{R}^{3}$ and $\Omega$, respectively. (In our convention, the mean curvature is the trace of the second fundamental form with respect to the outward unit normal.) Then we define the energy $E_{\lambda}$ as a direct generalization of the Liu–Yau energy [LY03, LY06]. It can be derived from the free-energy Hamiltonian in general relativity, see eq. (105) in [KIJ97].
Definition 2.1.
Let $(\Omega,g,k)$ be a compact initial data set with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature. Let $H$ be the mean curvature of $\Sigma$ in $\Omega$ with respect to the outward normal, and let $H_{0}$ be the mean curvature of the isometric embedding of $\Sigma$ into Euclidean space $\mathbb{R}^{3}$. Then
@@
As with the Liu–Yau energy, $E_{\lambda}$ can be expressed as the difference between the norms of the codimension-two mean curvature vectors in the relevant spacetimes. Indeed it’s not hard to see that
@@
Here $\vec{H}$ is the mean curvature vector of $\Sigma$ (assumed to be spacelike) within the original spacetime $(\mathscr{M},\mathscr{g})$ containing $(\Omega,g,k)$, while $\vec{H}_{0}$ is the mean curvature vector of $\Sigma$ within the de Sitter space $(\mathscr{M}_{0},\mathscr{g}_{0})$ with $\Lambda=3\lambda^{2}$ containing the compact initial data set $(\Omega_{0},\delta,k_{0}=\lambda\delta)$ within a constant $t$-slice. See figure 2 for a comparison between the Liu-Yau energy and $E_{\lambda}$. Like the Liu–Yau energy, inherent in the definition of $E_{\lambda}$, is the specific choice of Weyl embedding; different embeddings into de Sitter space will yield different energies. See section 7.1.
@@
Figure 2. Left: the Penrose diagram of Minkowski spacetime. To compute the Liu–Yau energy, the surface $\Sigma$ is isometrically embedded in a constant $t$-slice of Minkowski spacetime. Right: the Penrose diagram of de Sitter space. The surface is also isometrically embedded in a constant $t$-slice of the flat de Sitter space and lies within the cosmological horizon (shown in red).
@@ 3. Main conjecture and theorem
Consider an initial data set $(\Omega,g,k)$ where $\Omega$ is compact with non-empty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature. By the Weyl embedding theorem, which was solved independently by Nirenberg and Pogorelov [NIR53, POG52], $\Sigma$ can be isometrically embedded into the Euclidean space $\mathbb{R}^{3}$ with the image of $\Sigma$ as the boundary of a compact and convex domain $\Omega_{0}\subset\mathbb{R}^{3}$; moreover, this embedding is unique up to Euclidean rigid motions. Let $H_{0}$ be the mean curvature of the embedding in $\mathbb{R}^{3}$ with respect to the outward normal. In the initial data $\Omega$, let $H$ be the mean curvature of $\Sigma$ with respect to outward normal.
Our first conjecture states that the energy $E_{\lambda}$ should be nonnegative whenever the initial data set satisfies the dominant energy condition with respect to $\Lambda=3\lambda^{2}$ and the Weyl embedding of $\Sigma$ in $\mathbb{R}^{3}$ lies within the cosmological horizon; as in Liu-Yau [LY03, LY06], we also require that the mean curvature vectors of $\Sigma$ are spacelike within $\Omega$ and the Weyl embedding in $\mathbb{R}^{3}$ so that $E_{\lambda}$ is well defined.
@@
Fix $\lambda\geq 0$. Let $(\Omega,g,k)$ be a compact initial data set with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
@@
In the umbilic setting, $k=\lambda g$, the dominant energy condition in question is a purely Riemannian condition. Indeed, from the first constraint equation in (2), the dominant energy condition with $\Lambda=3\lambda^{2}$ is equivalent to requiring nonnegativity of scalar curvature, i.e. $R_{g}\geq 0$. This leads to the following purely Riemannian version of Conjecture 3.1.
@@
Fix $\lambda\geq 0$. Let $(\Omega,g,k=\lambda g)$ be a compact initial data set with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
@@
Addressing these conjectures involves two challenges. The first is geometric in nature: one would like to determine under which conditions the surface $\Sigma$ can be isometrically embedded inside the cosmological horizon. The second concerns the positivity of the energy $E_{\lambda}$ itself. We will show that both of these concerns can be resolved for some values of $\lambda$.
We start by introducing a constant $\alpha$ which depends on the initial data $(\Omega,g,k)$ via
$\alpha=\begin{cases}\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}&\text{if }\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\\ \min\left\{\sqrt{\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right)},\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}\right\}&\text{if }\sqrt{K_{\min}}\geq\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\end{cases}.$ (19)
In this expression, $K_{\min}$ is the minimum of the Gauss curvature $K$ on $\Sigma$, $r_{\Sigma}$ is the area radius of $\Sigma$, and $E_{\rm LY}$ is the Liu-Yau energy of the initial data (see equation (23) in the next section). We note that the definition of $\alpha$ does not involve $\lambda$ explicitly; however, an implicit dependence may occur if either the metric $g$ or the second fundamental form $k$ (and hence the Liu–Yau energy $E_{\rm LY}$) depends on $\lambda$.
@@
Fix $\lambda\geq 0$. Let $(\Omega,g,k)$ be a compact initial data set with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}$ holds. If $\lambda\leq\alpha$, where $\alpha$ is given by (19), then the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and the energy (15) is nonnegative:
@@
Moreover, if $0<\lambda<\alpha$, then $E_{\lambda}>0$; if $\lambda=\alpha>0$ and $E_{\lambda}=0$, then $\Sigma$ is isometric to a disjoint union of round spheres.
@@
    Fixed geometric data. One may fix an initial data set $(\Omega,g,k)$ that is completely independent of $\lambda$. In this case, $\lambda$ acts as an external parameter, and $\alpha$ is independent of $\lambda$. Therefore, in this case, the theorem provides an interval $0,\alpha]$ of admissible values of $\lambda$ for which the data embeds within the cosmological horizon and the quasi-local energy $E_{\lambda}$ is nonnegative.
@@
    $\lambda$-dependent data. Alternatively, one may consider a $\lambda$-dependent family of data $(\Omega,g(\lambda),k(\lambda))$, for instance determined by the constraint equations or by prescribing an umbilic second fundamental form. In this case, $\alpha$ depends on $\lambda$ since $E_{\rm LY}$ does. The theorem can be applied pointwise by verifying the condition $\lambda\leq\alpha$ for each value of $\lambda$.
Concerning the upper bound for $\lambda$ in Theorem 3.3 above, in our proof this appears as a technical restriction, which we believe to be non-optimal. This arises due to our appeals to Lemmas 5.1 and 5.2. Specifically, we appeal to Lemma 5.1 in order to grant the the Weyl embedding $\mathcal{W}:\Sigma\to\mathbb{R}^{3}$ isometrically embeds inside the cosmological horizon and for this we use a non-optimal estimate appealing to [JUN01] in order to produce a bound on the diameter of the enclosed region by $\mathcal{W}(\Sigma)$, which is suboptimal. This bound is what in the end sets an upper bound for $\lambda$ in the proof of Lemma 5.1. Similarly, we appeal to Lemma 5.2 to establish positive, where this lemma appeal to the stated upper bound for $\lambda$ to be able to reduce the positivity statement to that of the Liu–Yau energy $E_{\rm LY}$ (see Definition 4.2). Let us nevertheless highlight, that despite the introduction of such upper bounds for $\lambda$, Theorem 3.3 provides a geometrically determined such bound, establishing that for values of the cosmological constant which are not too large in comparison with clear-cut geometric quantities, the statement holds. Moreover, we emphasize that the actually observed value for the cosmological constant in our universe is remarkably small, and thus that, although maybe mathematically suboptimal, positive energy theorems for small $\Lambda$ are physically meaningful.